OpenStudy (anonymous):

can any one tell me how to graph this equation? -3x + 2y =8

OpenStudy (anonymous):

the easiest way to is get in into standard form. y=mx+b where m is slope and b is your intercept.... 2y=3x+8 y=3/2x+4

OpenStudy (anonymous):

start putting values in x and get corresponding value of y. plot y against x

OpenStudy (anonymous):

god your a life saver iv sat here for like ever searching it on google and no one helped -_- thank you @joshuamcelwain